Recent Log Entries

COMMENT ON Install new gun HV chamber and NEG/BPM beam pipe

Log Number: 3576248Logbook(s): Body: 

Helium leak test the system today, CEM mode on the RGA,
baseline in the 7e-10 Torr partial pressure helium. No leaks
detectable. System was warmed overnight to get a good
baseline for the leak test but will be kept at room
temperature for the weekend. Full bakeout of the system will
start Wednesday after power outage.

IOCSE14 off line

Log Number: 3576247Logbook(s): Body: 

I currently have IOCSE14 SEE bpm IOC off line. I found that the IOC had (2) bad cooling fans causing SEE bpm ADC cards to overheat. I have the fan tray in my office and currently I am looking for replacement fans. I also pulled the overheated ADC cards and installed spares.

gun leak check

Log Number: 3576246Logbook(s): Entry Makers: marcyBody: 

Careful leak check performed on Gun2 and NEG tube. CEM on for the RGA on the pump cart, no detectable leaks at all joints with helium and bag to accumulate.

The NEG tube and gun were warmed to ~70C overnight to get a good baseline for the leak test.
We'll have the system cool to room temperature over the weekend - we don't want the thin NEG coatings on the tube and gun to be pumping the gas in the system all weekend at elevated temperatures.

Tuesday we'll set up the bakeout, and Wednesday we'll start after the power outage in the injector.

COMMENT ON Transformer Electrical & Oil Testing

Log Number: 3576242Logbook(s): Body: 

I've attached three documents which provide a map of the
circuits that will be offline due to the work on Tuesday.
May 29th. Keep in mind, these circuit maps were produced by
software that is in development and, therefore, may not be
perfectly accurate yet.

iocse14 restored

Log Number: 3576238Logbook(s): Tags: Body: myRestore 2.12 Target restore date was 2018-05-25 08:19:12 The output from the myRestore execution was as follows: iocse14 system 2(2) restored seebpm 449(449) restored fastfb 44(44) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse14 (success)

COMMENT ON *General CAMAC / VME Crate Maintenance for 2018

Log Number: 3576236Logbook(s): Body: 

All circuit cards from CAMAC crate #2 have been cleaned
inspected and 1 was repaired. The cards were then
reinsstalled in the crate. With an assist from Scott Windham
wee removed the VME craate and took it to the TED for
cleaning , inspection. The power supplt was extremely dirty
but cleaned up nicely. We applied Duct Tape to an open slot
to prevent ingress of creatures and dirt since it is real
close to the floor.
After returning the VME crate to its location, I reenergized
the VME craate and both CAMAC Craates.

COMMENT ON *General CAMAC / VME Crate Maintenance for 2018

Log Number: 3576235Logbook(s): Body: 

All cards have been reinstalled in Crate #2. The VME crate
was pulled, with Scott Windham assist, cleaned, and
examined. It was extremely dirty. We added some Duct tape
across one open slot at the bottom to stop ingress of
creatures and dirt.
After all cards were returned to Crate #2 and the VME crate
was installed and hooked up I turned everything back on. The
2 CAMAC Crates came out of bypass and begun scanning as they
should. The serial highway leds look normal.
Only thing remaining is reinstallation of the front panel
connectors on the 2 CAMAC crates.

iocse14 restored

Log Number: 3576234Logbook(s): Tags: Body: myRestore 2.12 Target restore date was 2018-05-25 06:58:41 The output from the myRestore execution was as follows: iocse14 system 2(2) restored seebpm 449(449) restored fastfb 44(44) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse14 (success)

iocse9 restored

Log Number: 3576232Logbook(s): Tags: Body: myRestore 2.12 Target restore date was 2018-05-25 07:15:18 The output from the myRestore execution was as follows: iocse9 system 2(2) restored seebpm 449(449) restored fastfb 44(44) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse9 (success)

beam line vacuum around 1L07

Log Number: 3576231Logbook(s): Entry Makers: beneschBody: 

Figure 1 all. Figure 2 1L06 (not bad). Figure 3 1L08 probably needs 30K cycle. Figure 4, 1L09 ditto. Figure 5, 1L10, I don't understand why pressure rise lasted so long but it probably should be 30K cycled. Figure 6 1L11 OK.

iocse18 restored

Log Number: 3576229Logbook(s): Tags: Body: myRestore 2.12 Target restore date was 2018-05-25 07:01:51 The output from the myRestore execution was as follows: iocse18 system 2(2) restored seebpm 673(673) restored Execute: /usr/csite/certified/bin/ced2Epics -ced ced -A -s -i iocse18 (success)

COMMENT ON *General CAMAC / VME Crate Maintenance for 2018

Log Number: 3576226Logbook(s): Body: 

I inspected all of the cards pulled from crate #2 yesterday
afternoon. I found 1 bad switch. LT provided me with a
replacement switch yesterday afternoon for one of the Analog
Scanners.. I installed it this morning and I am beginning
the cleaning regiment on the circuit cards from Crate #2
now. Should be ready for reinsstallation by 8:00ish.

Pages